# Solving the Poisson equation using coupled Markov chains

### v3

@@ 2.3 Interpretation and verification of Assumption 2
Following Douc et al. (2018, Example 17.2.3), we may consider an independent Metropolis–Rosenbluth–Teller–Hastings kernel $P$ with target $\pi={\rm Uniform}(0,1)$ and proposal density $q(x)=(r+1)x^{r}$ on $(0,1)$, for some $r>0$. Since $\pi(x)/q(x)$ is unbounded, the resulting Markov chain cannot be geometrically ergodic (Mengersen & Tweedie, 1996) so Assumption 9 cannot hold. However, Assumption 12 does hold, as we now outline. Letting $s\in(\max(r,1),r+1)$ one can take $V(x)=x^{-s}$ and verify that
@@ 2.4 Illustration
We consider a target $\pi$ defined as the posterior distribution in a Cauchy location model with parameter $\theta$. We observe $n=3$ measurements $z_{1}=-8,z_{2}=+8,z_{3}=+17$, assumed to be realizations of $\text{Cauchy}(\theta,1)$. The prior on $\theta$ is $\text{Normal}(0,100)$. Since the prior is Normal and the likelihood is upper-bounded, $h:x\mapsto x$ is in $L^{m}(\pi)$ for all $m>0$. We consider the Gibbs sampler described in Robert (1995), that introduces auxiliary variables $\eta$ and alternates between Exponential and Normal draws as follows:
@@
where $\sigma^{2}=100$ is the prior variance. The coupling of this Gibbs sampler is done using common random numbers for the $\eta$-variables, and a maximal coupling for the update of $\theta$, as described in Appendix A.
We can verify Assumption 2. Consider any pair $\theta$, $\tilde{\theta}$, and the pair of next values, $\theta^{\prime}$, $\tilde{\theta}^{\prime}$. Write $\eta$ and $\tilde{\eta}$ for the auxiliary variables in each chain. First observe that the means of the Normal distributions, being of the form
@@
take values within $(-a,+a)$ with $a=\max|z_{i}|$, since they are weighted averages of $(z_{i})$ multiplied by a value in $(0,1)$. Therefore, the mean of the next $\theta^{\prime}$ is within a finite interval that does not depend on the previous $\theta$. Regarding the variance $(\sum_{i=1}^{n}\eta_{i}+\sigma^{-2})^{-1}$, note that $\eta_{i}=-{2}/({1+(\theta-z_{i})^{2}})\log U_{i}$, where $U_{i}$ is Uniform$(0,1)$, thus $\eta_{i}\leq-2\log U_{i}$ and finally $(\sum_{i=1}^{n}\eta_{i}+\sigma^{-2})^{-1}\geq(\sigma^{-2}+\sum_{i=1}^{n}(-2\log U_{i}))^{-1}$. Also, $(\sum_{i=1}^{n}\eta_{i}+\sigma^{-2})^{-1}\leq\sigma^{2}$. The distribution of $\sum_{i=1}^{n}(-2\log U_{i})$ does not depend on $\theta$, thus we can define an interval $(c,d)\subset(0,\infty)$ and $\epsilon\in(0,1)$, independently of $\theta$ and $\tilde{\theta}$, such that $(\sum_{i=1}^{n}\eta_{i}+\sigma^{-2})^{-1}\in(c,d)$ and $(\sum_{i=1}^{n}\tilde{\eta}_{i}+\sigma^{-2})^{-1}\in(c,d)$ simultaneously with probability $\epsilon$. Therefore, with probability $\epsilon$, the pair $(\theta^{\prime},\tilde{\theta}^{\prime})$ is drawn from a maximal coupling of two Normals, which means and variances are in finite intervals defined independently of $(\theta,\tilde{\theta})$. Two such Normals have a total variation distance that is bounded away from one, thus there exists some $\delta>0$ such that, overall, $\mathbb{P}(\theta^{\prime}=\tilde{\theta}^{\prime}|\theta,\tilde{\theta})>\epsilon\delta$, for some $\epsilon>0$, $\delta>0$ and all $(\theta$,$\tilde{\theta})$. Hence, Assumption 2 holds for all $\kappa\geq 1$.
The second algorithm is a Metropolis–Rosenbluth–Teller–Hastings (MRTH) algorithm with Normal proposal on $\theta$, with standard deviation $10$. Its coupling employs a reflection-maximal coupling of the proposals as described in Appendix A. Verification of Assumption 2 can be done as described in Section 2.3 via a drift condition (Jarner & Hansen, 2000), since the target is super-exponential. Thus, Assumption 2 holds again for all $\kappa\geq 1$.
@@
The state $y$ used to define $g=g_{y}$ in (2.3) is set to zero. Figure 2.1(a) shows the estimated fishy functions for the Gibbs sampler and MRTH, and Figure 2.1(b) shows the estimated second moments $\mathbb{E}[G(x)^{2}]$, for a grid of values of $x$ and using $10^{3}$ independent repeats of $G(x)$ for each $x$. According to Theorem 6, $\mathbb{E}[|G(x)|^{p}|<\infty$ for all $p\geq 1$, for both algorithms. Figure 2.1(a) shows that the fishy functions $g$ are markedly different for both algorithms. In particular, for the Gibbs sampler the fishy function appears to be uniformly bounded, whereas it appears to diverge for MRTH. To understand this difference, we recall an interpretation of fishy functions due to Kontoyiannis & Dellaportas (2009, Section 4). They observe that the average $t^{-1}\sum_{s=0}^{t-1}h(X_{s})$ has expectation $t^{-1}\sum_{s=0}^{t-1}P^{s}h(x)$ given $X_{0}=x$, thus $g_{\star}(x)$ in (1) represents the leading term in the asymptotic bias:
@@
Thus $g_{\star}(x)$ can be interpreted as the asymptotic influence of a starting point $X_{0}=x$ in the MCMC estimate. Figure 2.1(a) shows that this influence is unbounded for the random walk MRTH algorithm, which is intuitive: more MRTH steps are required for the chain to reach stationarity as $|X_{0}|$ increases. The same is not true for this particular Gibbs sampler. Indeed, an arbitrary large initial value for $\theta_{0}$ would result in $\eta$-variables arbitrarily close to zero, which in turn would result in $\theta_{1}$ being distributed approximately as the prior distribution $\text{Normal}(0,\sigma^{2})$. Thus, the chain returns to the center of the space in one step, from an arbitrary initial state.
@@ 3.1 Recovering Glynn–Rhee estimators from the Poisson equation
Glynn & Rhee (2014) show how certain couplings of Markov chains could be simulated so as to construct unbiased estimators of $\pi(h)$, for a class of test functions $h$, that can be computed for a random but finite cost. Under the assumption that the expectation of the cost and that the variance of the estimator are finite, then averages of independent copies converge at the Monte Carlo rate to $\pi(h)$ (Glynn & Whitt, 1992). Such estimators can be advantageous compared to ergodic average MCMC estimators, as independent copies can be produced by parallel machines. We refer to Section 3.3 in Jacob et al. (2020) for a short review on the use of unbiased estimators generated by parallel machines. In this section we explain how the unbiased estimators of Glynn & Rhee (2014), as well as the variants of Jacob et al. (2020), can be derived from the Poisson equation, and we provide novel results on the moments of these estimators.
@@
where the left-hand side does not depend on $x$ and $g$ is any fishy function for $h$. It then seems natural to estimate $\pi(h)$ by estimating the terms on the right-hand side, for any $x$. With $g=g_{\star}$ we may write
@@
where the right-hand side is a familiar quantity in the light of Glynn & Rhee (2014). In particular, one may run Algorithm 2.1 with $L=1$: starting from $X_{0}=Y_{0}=x$, sample $X_{1}\sim P(X_{0},\cdot)$, and iteratively sample $(X_{t+1},Y_{t})\sim\bar{P}((X_{t},Y_{t-1}),\cdot)$ for $t\geq 1$, where $\bar{P}$ is as defined in Section 2.1. The generated $(X,Y)$ process is such that $(X_{t+L},Y_{t})_{t\geq 0}$ is Markov and $X_{t+L}=Y_{t}$ almost surely for all $t$ large enough. Since $P^{t}h(x)$ is the expectation of both $h(X_{t})$ and $h(Y_{t})$, the estimator $h(X_{0})+\sum_{t=0}^{\infty}\{h(X_{t+1})-h(Y_{t})\}$ is unbiased for $\pi(h)$ under suitable assumptions.
@@
$=h(x)+\mathbb{E}_{x}\left[g_{\star}(X_{1})\right]-g_{\star}(x)$
$=h(x)+Pg_{\star}(x)-g_{\star}(x)$
$=\pi(h).$
The initialization of the chains can be generalized from a point mass at $x\in\mathbb{X}$ to any probability distribution $\pi_{0}$. Indeed, a re-arranged and integrated Poisson equation is
$\pi(h)=\pi_{0}(h)+\pi_{0}P(g)-\pi_{0}(g),$ (3.3)
@@
where marginally $X_{0}^{\prime}\sim\mu$, $Y_{0}^{\prime}\sim\mu$ and $X_{1}^{\prime}\sim\mu P$ for some probability measure $\mu$. We denote by $\gamma$ the joint probability measure for $(X_{1}^{\prime},Y_{0}^{\prime})$, since this features in our analysis, noting that this is a coupling of $\mu P$ and $\mu$.
The estimator in Definition 15 is identical to the estimator of Glynn & Rhee (2014), denoted by $H_{0}(X,Y)$ in Jacob et al. (2020), if one chooses $X_{1}^{\prime}\sim P(X_{0}^{\prime},\cdot)$ and $(X_{0}^{\prime},Y_{0}^{\prime})$ is drawn from some coupling of $\mu=\pi_{0}$ with itself.
@@ 3.2 Recovering unbiased MCMC estimators
We can also retrieve the more efficient variants proposed in Jacob et al. (2020). By changing the initial distribution $\gamma$, Definition 15 also admits the estimators denoted by $H_{k}(X,Y)$ for some $k\in\mathbb{N}$ in Jacob et al. (2020), and the estimators denoted by $H_{k:m}(X,Y)$ are obtained as averages of $H_{k}(X,Y)$ over a range of values of $k$. Unbiased estimators based on chains coupled with a lag $L>1$ (Vanetti & Doucet, 2020) can be retrieved as well by considering the Poisson equation associated with the iterated Markov kernel $P^{L}$. To make this precise, we make the following definition.
@@
$H_{k}:=h(X_{k}^{\prime})+\sum_{j=1}^{\infty}h(X_{k+jL}^{\prime})-h(Y_{k+(j-1)L}^{\prime}),$
where $k\in\mathbb{N}$, $L\geq 1$, $(X_{t+L}^{\prime},Y_{t}^{\prime})_{t\geq 0}$ is a time-homogeneous Markov chain with Markov kernel $\bar{P}$, and $(X_{t}^{\prime})_{t=0}^{L}$ is a Markov chain with transition kernel $P$ and initial distribution $\pi_{0}$ and $Y^{\prime}_{0}\sim\pi_{0}$ independent of $(X^{\prime}_{0:L})$. In particular, $(X_{L}^{\prime},Y^{\prime}_{0})\sim\gamma=\pi_{0}P^{L}\otimes\pi_{0}$. The coupled chain may be generated with Algorithm 2.1, starting from two independent draws from $\pi_{0}$, and the meeting time $\tau$ is here $\inf\{t>L:X_{t}=Y_{t-L}\}$. We define for any $k,\ell\in\mathbb{N}$ with $k\leq\ell$, the average of such estimators as
$H_{k:\ell}:=\frac{1}{\ell-k+1}\sum_{t=k}^{\ell}H_{t}.$
It will be convenient to view the unbiased MCMC estimator $H_{k:\ell}$ in Definition 16 as equivalent to the expectation of $h$ with respect to an unbiased MCMC signed measure $\hat{\pi}_{k:\ell}$. After some routine calculations described in Appendix C, and replacing evaluations of the test function $h$ by Dirac delta masses, one obtains the following signed measure
@@
with
$v_{t}=\lfloor(t-k)/L\rfloor-\lceil\max(L,t-\ell)/L\rceil+1.$ (3.6)
In the following developments, we will write $\hat{\pi}$ instead of $\hat{\pi}_{k:\ell}$, and we may write $\hat{\pi}=\sum_{i=1}^{N}\omega_{i}\delta_{Z_{i}}$, instead of the full expression (3.5). For such measures and a function $h$, $\hat{\pi}(h)=\sum_{i=1}^{N}\omega_{i}h(Z_{i})=H_{k:\ell}$.
The following result is established in Appendix B.4, as a particular case of Proposition 43, where upper bounds are given for $\mathbb{E}[|H_{k:\ell}|^{p}]$. This result can be compared with Jacob et al. (Proposition 1, 2020) and Middleton et al. (Theorem 1, 2020), which provide only finite second moments. The latter obtains the same conditions on $\kappa$ and $m$ for $p=2$. The bounded ${\rm d}\pi_{0}/{\rm d}\pi$ assumption allows one to avoid the less explicit assumption that $\sup_{n\geq 0}\mu P^{n}(|h|^{2+\eta})<\infty$ and can often be verified in practice.
@@
Under Assumption 2, let $h\in L^{m}(\pi)$ with $m>\kappa/(\kappa-1)$, and ${\rm d}\pi_{0}/{\rm d}\pi\leq M$ with $M<\infty$. Then for any $k,\ell\in\mathbb{N}$ with $k\leq\ell$, any $L\geq 1$, $\mathbb{E}[H_{k:\ell}]=\pi(h)$ and for $p\geq 1$ such that $\frac{1}{p}>\frac{1}{m}+\frac{1}{\kappa}$, $\mathbb{E}[|H_{k:\ell}|^{p}]<\infty$.
@@
Assume the conditions of Theorem 3 and let $k\in\mathbb{N}$, and let $L\geq 1$. Then $H_{k:\ell}$ in Definition 16 satisfies
@@
as $\ell\to\infty$.
Proof.
As recalled in Appendix C, we may write $H_{t}=h(X_{t}^{\prime})+B_{t}$. It follows that we may write
$\sqrt{\ell-k+1}\left(H_{k:\ell}-\pi(h)\right)=\frac{1}{\sqrt{\ell-k+1}}\sum_{t=k}^{\ell}h_{0}(X_{t}^{\prime})+\frac{1}{\sqrt{\ell-k+1}}\sum_{t=k}^{\ell}B_{t}.$
Under these conditions, for suitably large $\ell-k$ the concentration of $H_{k:\ell}$ is similar to that of the ergodic average MCMC estimator of a similar computational cost, noting that one is simulating only a single chain after the meeting time. Proposition 19 therefore suggests that the efficiency of unbiased MCMC estimators can be made arbitrarily close to that of ergodic average MCMC, by choosing a large enough $\ell$. For a fixed choice of $\ell$, one can assess the efficiency loss by comparing the product of expected cost and variance of unbiased MCMC, with the asymptotic variance $v(P,h)$. However, the numerical approximation of the asymptotic variance $v(P,h)$ is not easy, and typically requires long chains. Accordingly, Glynn & Rhee (2014); Agapiou et al. (2018); Jacob et al. (2020) all compare the efficiency of unbiased estimators with that of ergodic averages obtained from long runs. In Section 4 we propose a new asymptotic variance estimator that is itself unbiased, and thus can be generated independently on parallel machines, for a random but finite time, and averaged to obtain a consistent estimator of $v(P,h)$ converging at the Monte Carlo rate. This alleviates the need for long chains, not only for the estimation of $\pi(h)$ but also for the efficiency assessment of unbiased MCMC.

### v4

@@ 2.3 Interpretation and verification of Assumption 2
The above assumptions and proposition can be compared to Assumptions 4–5 and Theorem 2 in Middleton et al. (2020). Our assumptions are similar but slightly weaker, namely we do not assume that $P$ is aperiodic since it is implied by the other conditions as stated in Proposition B.3. If we were to employ Theorem 2 in Middleton et al. (2020) to directly verify Assumption 2, setting $\pi_{0}$ to $\pi$, we would have to assume that $\pi$ is supported on a compact set. This is the main motivation for our Proposition 13. On the other hand, our result establishes less than $\alpha/(1-\alpha)$ moments of $\tau$, whereas Middleton et al. (Theorem 2, 2020) obtain $1/(1-\alpha)$ moments, which is strictly larger.
Proposition 13 is useful if a polynomial drift condition holds for the MCMC algorithm under consideration. Andrieu & Vihola (2015) obtain such drift conditions for pseudo-marginal methods, such as random walk MRTH where target density evaluations are replaced by unbiased estimates, see their Corollary 31, Theorem 38 and Theorem 45. Using these results, Middleton et al. (Section 2.3, 2020) proceed to verify the assumptions of their Theorem 2 in the context of pseudo-marginal MRTH, under conditions on the noise of the target density estimators and the tails of $\pi$. Under these conditions stated in their Proposition 2, our assumptions are also satisfied. In the case of MRTH with independent proposals, we provide details on the verification of the assumptions of Proposition 13 in the following toy example.
@@
Following Douc et al. (2018, Example 17.2.3), we consider an independent MRTH kernel $P$ with target $\pi={\rm Uniform}(0,1)$ and proposal density $q(x)=(r+1)x^{r}$ on $(0,1)$, for some $r>0$. Since $\pi(x)/q(x)$ is unbounded, the resulting Markov chain cannot be geometrically ergodic (Mengersen & Tweedie, 1996) so Assumption 9 cannot hold. However, Assumption 12 does hold. Indeed, letting $s\in(\max(r,1),r+1)$ one can take $V(x)=x^{-s}$ and verify that
@@
The tails of the meeting times can also be studied directly, without drift conditions, as is done in Lee et al. (2020); Karjalainen et al. (2023) for couplings of conditional particle filters, in Nguyen et al. (2022) for couplings of a Gibbs sampler for Bayesian clustering with Dirichlet process mixtures, in Deligiannidis et al. (2024) for the common random numbers coupling of MRTH with independent proposals, and in Section 2.4 below.
@@ 2.4 Illustration
We illustrate the fishy function estimator of Definition 5 and the verification of Assumption 2 in the following example. The target $\pi$ is defined as the posterior in a Cauchy location model with parameter $\theta$. We observe $n=3$ measurements $z_{1}=-8,z_{2}=+8,z_{3}=+17$, assumed to be realizations of $\text{Cauchy}(\theta,1)$. The prior on $\theta$ is $\text{Normal}(0,100)$. Since the prior is Normal and the likelihood is upper-bounded, $h:x\mapsto x$ is in $L^{m}(\pi)$ for all $m>0$. We consider the Gibbs sampler described in Robert (1995), that introduces auxiliary variables $\eta$ and alternates between Exponential and Normal draws as follows:
@@
where $\sigma^{2}=100$ is the prior variance. The coupling of this Gibbs sampler is done using common random numbers for the $\eta$-variables, and a maximal coupling for the update of $\theta$.
We can verify Assumption 2 by direct calculations. Consider any pair $\theta$, $\tilde{\theta}$, and the pair of next values, $\theta^{\prime}$, $\tilde{\theta}^{\prime}$. Write $\eta$ and $\tilde{\eta}$ for the auxiliary variables in each chain. First observe that the means of the Normal distributions, being of the form
@@
take values within $(-a,+a)$ with $a=\max|z_{i}|$, since they are weighted averages of $(z_{i})$ multiplied by a value in $(0,1)$. Therefore, the mean of the next $\theta^{\prime}$ is within a finite interval that does not depend on the previous $\theta$. Regarding the variance $(\sum_{i=1}^{n}\eta_{i}+\sigma^{-2})^{-1}$, note that $\eta_{i}=-{2}/({1+(\theta-z_{i})^{2}})\log U_{i}$, where $U_{i}$ is Uniform$(0,1)$, thus $\eta_{i}\leq-2\log U_{i}$ and finally $(\sum_{i=1}^{n}\eta_{i}+\sigma^{-2})^{-1}\geq(\sigma^{-2}+\sum_{i=1}^{n}(-2\log U_{i}))^{-1}$. Also, $(\sum_{i=1}^{n}\eta_{i}+\sigma^{-2})^{-1}\leq\sigma^{2}$. The distribution of $\sum_{i=1}^{n}(-2\log U_{i})$ does not depend on $\theta$, thus we can define an interval $(c,d)\subset(0,\infty)$ and $\epsilon\in(0,1)$, independently of $\theta$ and $\tilde{\theta}$, such that $(\sum_{i=1}^{n}\eta_{i}+\sigma^{-2})^{-1}\in(c,d)$ and $(\sum_{i=1}^{n}\tilde{\eta}_{i}+\sigma^{-2})^{-1}\in(c,d)$ simultaneously with probability $\epsilon$. Therefore, with probability $\epsilon$, $(\theta^{\prime},\tilde{\theta}^{\prime})$ is drawn from a maximal coupling of two Normals, which means and variances are in finite intervals defined independently of $(\theta,\tilde{\theta})$. Two such Normals have a total variation distance that is bounded away from one, thus there exists $\delta>0$ such that, $\mathbb{P}(\theta^{\prime}=\tilde{\theta}^{\prime}|\theta,\tilde{\theta})>\epsilon\delta$, for some $\epsilon>0$, $\delta>0$ and all $(\theta$,$\tilde{\theta})$. Hence, Assumption 2 holds for all $\kappa\geq 1$.
The second algorithm is a Metropolis–Rosenbluth–Teller–Hastings (MRTH) algorithm with Normal proposal on $\theta$, with standard deviation $10$. Its coupling employs a reflection-maximal coupling of the proposals as described in Appendix A.1. Verification of Assumption 2 can be done as described in Section 2.3 via a drift condition (Jarner & Hansen, 2000), since the target is super-exponential. Thus, Assumption 2 holds again for all $\kappa\geq 1$.
@@
The state $y$ used to define $g=g_{y}$ in (2.3) is set to zero. Figure 2.1(a) shows the estimated fishy functions for the Gibbs sampler and MRTH, and Figure 2.1(b) shows the estimated second moments $\mathbb{E}[G(x)^{2}]$, for a grid of values of $x$ and using $10^{3}$ independent repeats of $G(x)$ for each $x$. According to Theorem 6, $\mathbb{E}[|G(x)|^{p}|<\infty$ for all $p\geq 1$, for both algorithms. Figure 2.1(a) shows that the fishy functions $g$ are markedly different for both algorithms. For the Gibbs sampler the fishy function appears uniformly bounded, whereas it seems to diverge for MRTH. To understand this difference, we recall an interpretation of fishy functions due to Kontoyiannis & Dellaportas (2009, Section 4). The average $t^{-1}\sum_{s=0}^{t-1}h(X_{s})$ has expectation $t^{-1}\sum_{s=0}^{t-1}P^{s}h(x)$ given $X_{0}=x$, thus $g_{\star}(x)$ in (1) represents the leading term in the asymptotic bias:
@@
Thus $g_{\star}(x)$ measures the asymptotic influence of a starting point $X_{0}=x$ in the MCMC estimate. Figure 2.1(a) shows that this influence is unbounded for the random walk MRTH algorithm, which is intuitive: more MRTH steps are required for the chain to reach stationarity as $|X_{0}|$ increases. The same is not true for the Gibbs sampler. Indeed, an arbitrary large initial value for $\theta_{0}$ would result in $\eta$-variables arbitrarily close to zero, which in turn would result in $\theta_{1}$ being distributed approximately as the prior distribution $\text{Normal}(0,\sigma^{2})$. Thus, the chain returns to the center of the space in one step, from an arbitrary initial state.
@@ 3.1 Recovering Glynn–Rhee estimators from the Poisson equation
Glynn & Rhee (2014) show how certain couplings of Markov chains could be simulated so as to construct unbiased estimators of $\pi(h)$, for a class of test functions $h$, that can be computed for a random but finite cost. If the expected cost and the variance of an estimator are finite, then averages of independent copies converge at the Monte Carlo rate to $\pi(h)$ (Glynn & Whitt, 1992). Such estimators can be advantageous compared to ergodic average MCMC estimators, as independent copies can be produced by parallel machines. Section 3.3 in Jacob et al. (2020) provides a review on the use of unbiased estimators generated by parallel machines. In this section we explain how the unbiased estimators of Glynn & Rhee (2014), as well as the variants of Jacob et al. (2020), can be derived from the Poisson equation, and we provide conditions implying the finiteness on their moments of any order (Theorem 17). In Proposition 19 we provide an asymptotic result facilitating the comparison between standard and unbiased MCMC estimators. Finally, in Section 3.3 we discuss the equivalent of “thinning” in unbiased MCMC, which is useful for the developments of Section 4.
@@
where the left-hand side does not depend on $x$ and $g$ is any fishy function for $h$. It then seems natural to estimate $\pi(h)$ by estimating the terms on the right-hand side, for any $x$. With $g=g_{\star}$ as in Definition 1 we may write
@@
where the right-hand side is a familiar quantity in the light of Glynn & Rhee (2014). In particular, one may run Algorithm 2.1 with $L=1$: starting from $X_{0}=Y_{0}=x$, sample $X_{1}\sim P(X_{0},\cdot)$, and iteratively sample $(X_{t+1},Y_{t})$ using the transition $\bar{P}$ defined in Section 2.1. The generated $(X,Y)$ process is such that $X_{t+1}=Y_{t}$ almost surely for all $t$ large enough. Since $P^{t}h(x)$ is the expectation of both $h(X_{t})$ and $h(Y_{t})$, the estimator $h(X_{0})+\sum_{t=0}^{\infty}\{h(X_{t+1})-h(Y_{t})\}$ is unbiased for $\pi(h)$ under suitable assumptions (see Theorem 17).
@@
$=h(x)+\mathbb{E}_{x}\left[g_{\star}(X_{1})\right]-g_{\star}(x)=\pi(h).$
The initialization of the chains can be generalized from a point mass at $x\in\mathbb{X}$ to a probability distribution $\mu$ (a condition on the initial distribution will be stated in Theorem 17). Indeed, a re-arranged and integrated Poisson equation is
$\pi(h)=\mu(h)+\mu P(g)-\mu(g),$ (3.3)
@@
where marginally $X_{0}^{\prime}\sim\mu$, $Y_{0}^{\prime}\sim\mu$ and $X_{1}^{\prime}\sim\mu P$ for some probability measure $\mu$, and $G$ is in Definition 5. We denote by $\gamma$ the joint probability measure for $(X_{1}^{\prime},Y_{0}^{\prime})$, since this features in our analysis, noting that this is a coupling of $\mu P$ and $\mu$.
The estimator in Definition 15 is identical to the estimator of Glynn & Rhee (2014), denoted by $H_{0}(X,Y)$ in Jacob et al. (2020), if one chooses $X_{1}^{\prime}\sim P(X_{0}^{\prime},\cdot)$ and $(X_{0}^{\prime},Y_{0}^{\prime})$ is drawn from some coupling of $\mu=\pi_{0}$ with itself. In the following we focus on an independent initialization of the chains.
@@ 3.2 Recovering unbiased MCMC estimators
We can also retrieve the more efficient variants proposed in Jacob et al. (2020). By changing the initial distribution $\gamma$, Definition 15 also admits the estimators denoted by $H_{k}(X,Y)$ for some $k\in\mathbb{N}$ in Jacob et al. (2020), and the estimators denoted by $H_{k:m}(X,Y)$ are obtained as averages of $H_{k}(X,Y)$ over a range of values of $k$. Unbiased estimators based on chains coupled with a lag $L>1$ (Vanetti & Doucet, 2020) can be retrieved as well by considering the Poisson equation associated with the iterated Markov kernel $P^{L}$. To make this precise, we present the following definition.
@@
$H_{k}:=h(X_{k})+\sum_{j=1}^{\infty}h(X_{k+jL})-h(Y_{k+(j-1)L}),$
where $k\in\mathbb{N}$, $L\geq 1$, $(X_{t+L},Y_{t})_{t\geq 0}$ is a time-homogeneous Markov chain with Markov kernel $\bar{P}$, and $(X_{t})_{t=0}^{L}$ is a Markov chain with transition kernel $P$ and initial distribution $\pi_{0}$ and $Y_{0}\sim\pi_{0}$ independent of $(X_{0},\ldots,X_{L})$. In particular, $(X_{L},Y_{0})\sim\pi_{0}P^{L}\otimes\pi_{0}$. The coupled chain may be generated with Algorithm 2.1, starting from two independent draws from $\pi_{0}$, and the meeting time $\tau$ here is $\inf\{t>L:X_{t}=Y_{t-L}\}$. We define for any $k,\ell\in\mathbb{N}$ with $k\leq\ell$, the average of such estimators as $H_{k:\ell}:={(\ell-k+1)}^{-1}\sum_{t=k}^{\ell}H_{t}$.
It will be convenient to view the unbiased MCMC estimator $H_{k:\ell}$ in Definition 16 as equivalent to the expectation of $h$ with respect to an unbiased MCMC signed measure $\hat{\pi}_{k:\ell}$. After some routine calculations described in Appendix A.2, and replacing evaluations of the test function $h$ by Dirac delta masses, one obtains the following signed measure
@@
$\text{with}\quad v_{t}$ $=\lfloor(t-k)/L\rfloor-\lceil\max(L,t-\ell)/L\rceil+1.$ (3.6)
The following result is established in Appendix B.4, as a particular case of Proposition B.20, where upper bounds are given for $\mathbb{E}[|H_{k:\ell}|^{p}]$. This result can be compared with Jacob et al. (Proposition 1, 2020) and Middleton et al. (Theorem 1, 2020), which provide only finite second moments. The latter obtains the same conditions on $\kappa$ and $m$ for $p=2$. The bounded ${\rm d}\pi_{0}/{\rm d}\pi$ assumption allows one to avoid the less explicit assumption that $\sup_{n\geq 0}\pi_{0}P^{n}(|h|^{2+\eta})<\infty$ and can often be verified as the MCMC user chooses $\pi_{0}$.
@@
Under Assumption 2, let $h\in L^{m}(\pi)$ with $m>\kappa/(\kappa-1)$, and ${\rm d}\pi_{0}/{\rm d}\pi\leq M$ with $M<\infty$. Then, with $H_{k:\ell}$ in Definition 16 for any $k,\ell\in\mathbb{N}$ with $k\leq\ell$, any $L\geq 1$, $\mathbb{E}[H_{k:\ell}]=\pi(h)$ and for $p\geq 1$ such that $\frac{1}{p}>\frac{1}{m}+\frac{1}{\kappa}$, $\mathbb{E}[|H_{k:\ell}|^{p}]<\infty$.
@@
Assume the conditions of Theorem 3 and let $k\in\mathbb{N}$, and let $L\geq 1$. Then $H_{k:\ell}$ in Definition 16 satisfies, as $\ell\to\infty$,
@@
Proposition 19 is established in Appendix B.4. Under the stated conditions, for suitably large $\ell-k$ the concentration of $H_{k:\ell}$ is similar to that of the ergodic average MCMC estimator of a similar computational cost, noting that one is simulating only a single chain after the meeting time. Proposition 19 suggests that the efficiency of unbiased MCMC estimators can be made arbitrarily close to that of ergodic average MCMC, by choosing a large enough $\ell$. For a fixed $\ell$, one can assess the efficiency loss by comparing the product of expected cost and variance of unbiased MCMC, with the asymptotic variance $v(P,h)$. However, the numerical approximation of $v(P,h)$ is not easy, and typically requires long chains. Accordingly, Glynn & Rhee (2014); Agapiou et al. (2018); Jacob et al. (2020) all compare the efficiency of unbiased estimators with that of ergodic averages obtained from long runs. In Section 4 we propose a new asymptotic variance estimator that is itself unbiased, and thus can be generated independently on parallel machines, for a random but finite time, and averaged to obtain a consistent estimator of $v(P,h)$ converging at the Monte Carlo rate. This alleviates the need for long chains, not only for the estimation of $\pi(h)$ but also for the efficiency assessment of unbiased MCMC.